# Fast Ewald Summation using Prolate Spheroidal Wave Functions

### v1

@@ 3.2 Fast Ewald summation
The steps of the fast Ewald method, as presented above, are summarized in Algorithm 2.
@@ 4 Fast Ewald summation for the Laplace kernel using the first PSWF of order zero
In this section, we present how the first PSWF of order zero, $\psi_{0}^{c}$, can be employed both as a mollifier in the Ewald-split and as a window function for the spreading/interpolation steps of Algorithm 2.
4.1 General Ewald decomposition for numerical mollifiers with support [1,1][-1,1]
@@
Finally, substituting (49) in the Ewald decomposition $G=M+R$ the residual kernel becomes
@@
4.2 A PSWF mollifier for the Ewald split
Following Section 4.1, we now construct a PSWF mollifier. We start by defining an unscaled mollifier $\gamma^{c_{s}}_{1}$, supported on [1,1][-1,1], as
$\gamma_{1}^{c_{s}}(x):=\begin{cases}\dfrac{\psi_{0}^{c_{s}}(x)}{\hat{\psi}_{0}^{c_{s}}(0)}=\dfrac{\psi_{0}^{c_{s}}(x)}{\lambda_{0}\psi_{0}^{c_{s}}(0)},&x\in[-1,1],\\[5.69054pt] 0,&x\notin[-1,1],\end{cases}$ (53)
where $\psi_{0}^{c_{s}}$ denotes the first PSWF with Fourier bandlimit $c_{s}>0$ (the subscript “s” stands for “split”). Here, $\lambda_{0}$ is the eigenvalue associated with $\psi_{0}^{c_{s}}$, and the identity $\hat{\psi}_{0}^{c_{s}}(0)=\lambda_{0}\psi_{0}^{c_{s}}(0)$ follows from Lemma 2. This normalization ensures that $\int_{\mathbb{R}}\gamma^{c_{s}}_{1}(x)\,\mathrm{d}x=\int_{-1}^{1}\gamma^{c_{s}}_{1}(x)\,\mathrm{d}x=1$. Then, to obtain a mollifier supported on a general radius $r_{c}>0$, we scale $\gamma^{c_{s}}_{1}$ as
@@ 4.2 A PSWF mollifier for the Ewald split
The Fourier transform of $\gamma_{r_{c}}^{c_{s}}$ admits a closed-form expression within the band $c_{s}/r_{c},\,c_{s}/r_{c}]$:
$\hat{\gamma}_{r_{c}}^{c_{s}}(\omega)=\hat{\gamma}_{1}^{c_{s}}(r_{c}\omega)=\frac{\psi_{0}^{c_{s}}\!\left(r_{c}\omega/c_{s}\right)}{\psi_{0}^{c_{s}}(0)},\qquad\omega\in\bigg[-\frac{c_{s}}{r_{c}},\,\frac{c_{s}}{r_{c}}\bigg].$ (55)
Outside this interval, the Fourier transform must be evaluated directly from its defining integral,
$\hat{\gamma}_{r_{c}}^{c_{s}}(\omega)=\int_{-r_{c}}^{r_{c}}\gamma_{r_{c}}^{c_{s}}(x)\,e^{-\mathrm{i}\omega x}\,\mathrm{d}x.$ (56)
We remark that (56) is included only for completeness; the integral representation (56) is never used in practice, since only Fourier modes within the band $c_{s}/r_{c},\,c_{s}/r_{c}]$ are required.
@@
$\Phi^{c_{s}}_{r_{c}}(x):=2\int_{0}^{x}\gamma^{c_{s}}_{r_{c}}(u)\,\mathrm{d}u=\frac{2}{\lambda_{0}\,\psi_{0}^{c_{s}}(0)}\int_{0}^{x/r_{c}}\psi_{0}^{c_{s}}(v)\,\mathrm{d}v,\qquad x\in[0,r_{c}],$ (57)
@@
$\hat{M}(\bm{\omega})=\frac{4\pi}{|\bm{\omega}|^{2}}\,\hat{\gamma}_{r_{c}}^{c_{s}}(|\bm{\omega}|)=\frac{4\pi}{|\bm{\omega}|^{2}}\frac{\psi_{0}^{c_{s}}\!\left(r_{c}|\bm{\omega}|/c_{s}\right)}{\psi_{0}^{c_{s}}(0)},\qquad|\bm{\omega}|\leq\frac{c_{s}}{r_{c}},$ (58)
@@
Table 1: Gaussian and PSWF-based mollifiers. The Gaussian mollifier has width parameter $\sigma$, while the PSWF mollifier is compactly supported with cutoff $r_{c}$ and bandlimit $c_{s}$. The fact that $\hat{\gamma}_{r_{c}}^{\mathrm{PSWF}}(\omega)$ does not admit a closed-form expression outside the interval $c_{s}/r_{c},\,c_{s}/r_{c}]$ is unimportant in practice, since this region is never used.
@@
$\gamma_{r_{c}}^{\text{PSWF}}(x)=\begin{cases}\dfrac{\psi_{0}^{c_{s}}(x/r_{c})}{r_{c}\,\lambda_{0}\,\psi_{0}^{c_{s}}(0)},&x\in[-r_{c},r_{c}]\\[5.81247pt] 0,&\text{otherwise}\end{cases}$ $\hat{\gamma}_{r_{c}}^{\text{PSWF}}(\omega)=\begin{cases}\displaystyle\frac{\psi_{0}^{c_{s}}\!\left(r_{c}\omega/c_{s}\right)}{\psi_{0}^{c_{s}}(0)}&\omega\in\left[-\frac{c_{s}}{r_{c}},\frac{c_{s}}{r_{c}}\right]\\[5.81247pt] \text{no closed form},&\text{otherwise}\end{cases}$
@@ 4.3 A PSWF window function
We define a one-dimensional, unscaled window function by
$\varphi_{1}^{c_{w}}(x):=\frac{\psi_{0}^{c_{w}}(x)}{\psi_{0}^{c_{w}}(0)},\qquad x\in\mathbb{R}.$ (59)
By Lemma 2, its Fourier transform for $\omega\in[-c_{w},c_{w}]$ is given by
@@
$\varphi^{c_{w}}(\bm{x}):=\prod_{i=1}^{3}\varphi_{1}^{c_{w}}(x_{i}/\alpha)=\prod_{i=1}^{3}\frac{\psi_{0}^{c_{w}}(x_{i}/\alpha)}{\psi_{0}^{c_{w}}(0)}.$ (61)
The Fourier transform of $\varphi^{c_{w}}$ has per-coordinate bandlimit $|\omega_{i}|\leq c_{w}/\alpha$ and is given by
@@
Finally, recalling Section 2.4, the truncated window is defined as
$\varphi_{t}^{c_{w}}(\bm{x}):=\begin{cases}\varphi^{c_{w}}(\bm{x}),&\bm{x}\in[-\alpha,\alpha]^{3},\\ 0,&\bm{x}\notin[-\alpha,\alpha]^{3},\end{cases}$ (63)
with periodization given by $\tilde{\varphi}_{t}^{c_{w}}(\bm{x})=\sum_{\bm{r}\in\mathbb{Z}^{3}}\varphi^{c_{w}}_{t}(\bm{x}+L\bm{r})$.
The truncated window $\varphi_{t}^{c_{w}}(\bm{x})$ with Fourier transform given by $\hat{\varphi}^{c_{w}}(\bm{\omega})$ in $c_{w}/\alpha,c_{w}/\alpha]^{3}$ is what we will refer to as the PSWF window function; see Table 2.
@@
Thus, for force computations (e.g. $\bm{F}_{i}=\rho_{i}\nabla\phi_{i}$), $\nabla\phi_{i}$ can be obtained by replacing the last interpolation step (step 5) in Algorithm 2 with
@@ 4.3.1 Comparison of the PSWF window with Gaussian and B-spline windows
$\varphi^{\text{PSWF}}(x)=\begin{cases}\dfrac{\psi_{0}^{c_{w}}(x/\alpha)}{\psi_{0}^{c_{w}}(0)},&x\in[-\alpha,\alpha]\\[5.81247pt] 0,&\text{otherwise}\end{cases}$ $\hat{\varphi}^{\text{PSWF}}(\omega)=\begin{cases}\displaystyle\alpha\,\lambda_{0}\,\frac{\psi_{0}^{c_{w}}(\alpha\omega/c_{w})}{\psi_{0}^{c_{w}}(0)},&\omega\in\Big[-\frac{c_{w}}{\alpha},\frac{c_{w}}{\alpha}\Big]\\[5.81247pt] \text{no closed form},&\text{otherwise}\end{cases}$
@@
B-splines have support width $P$, that equals the spline order, and have Fourier transform $\hat{\varphi}(\omega)=h\,\hat{B}_{P}(h\omega)$, where $h$ denotes the grid spacing and $\hat{B}_{P}(\omega)=(\sin(\omega/2)/(\omega/2))^{P}$. Consequently, their spectral decay is algebraic. For a fixed real-space width $\alpha$, B-splines therefore require either a finer grid or a higher order $P$ to match the accuracy of the PSWF window; increasing $P$ enlarges the support and increases the cost of spreading and interpolation. The widely used SPME method [11]—implemented, for example, in GROMACS [34]—employs non-centered cardinal B-splines for spreading and interpolation and deconvolves their spectrum through a modified diagonal scaling (see Remark 5). However, the underlying algebraic spectral decay remains unchanged.
@@ 5 Theoretical error analysis for the PSWF split and window functions
In this section, we present a rigorous theoretical analysis of the error sources arising in fast Ewald summation using PSWF-based functions. We derive absolute $L^{2}$ error estimates associated with both the PSWF mollifier and the PSWF window function introduced in Section 4, establishing exponential error decay and providing sharp bounds. For the purpose of parameter selection—where fully rigorous bounds are unnecessarily detailed—we also derive simplified, closed-form error models that are well suited for practical use.
@@
Let $\phi^{\text{far}}_{t}$ denote the truncated far-field sum $\phi^{\text{far}}$, and let $\phi^{\text{far}}_{t,h}$ denote an approximation of $\phi^{\text{far}}_{t}$ computed using the fast Ewald algorithm (Algorithm 2). In fast Ewald summation, two principal sources of error arise: (i) truncation of Fourier modes in the Fourier-transformed mollified kernel—referred to as split errors—quantified by $\frac{1}{\sqrt{V}}\|\phi^{\text{far}}-\phi_{t}^{\text{far}}\|_{L^{2}(\Omega)}$, and (ii) approximation errors arising in the spreading and interpolation steps, quantified by $\frac{1}{\sqrt{V}}\|\phi_{t}^{\text{far}}-\phi_{t,h}^{\text{far}}\|_{L^{2}(\Omega)}$.
@@
5.1 Ewald-split truncation errors
@@
The constant $C_{\rho}$ reflects the degree of cancellation among the oscillatory exponential terms. In the worst case (highly clustered particles), one may have $C_{\rho}$ as large as $n$. However, for well-distributed configurations (e.g., approximately uniform particle distributions), the exponential vectors corresponding to distinct wavenumbers are nearly orthogonal, leading to substantial cancellations. In such cases, it is reasonable to assume that $C_{\rho}\approx 1$.
We are now ready to state the following lemma, which provides an error estimate for the far-field direct sum for a general mollifier $\gamma_{\sigma}$.
Lemma 4.
Let $\phi_{t}^{\mathrm{far}}$ denote the Fourier-space far-field sum truncated to modes $|k_{i}|\leq K_{\text{max}}$ for $i=1,2,3$. Then the squared RMS-error of the truncation satisfies
$\frac{1}{V}\|\phi^{\mathrm{far}}-\phi_{t}^{\mathrm{far}}\|_{L^{2}(\Omega)}^{2}\leq\frac{8\,C_{\rho}\,\|\bm{\rho}\|_{2}^{2}}{V}\int_{\omega_{*}}^{\infty}|\hat{\gamma}_{\sigma}(\omega)|^{2}\frac{1}{\omega^{2}}\,d\omega,$ (72)
where $V=L^{3}$, $C_{\rho}$ is the constant from Lemma 3, and $\omega_{*}:=\max\!\big\{\omega_{\max}-\tfrac{\sqrt{3}\pi}{L},\,0\big\}$.
@@
$\frac{1}{V}\|\phi^{\mathrm{far}}-\phi_{t}^{\mathrm{far}}\|_{L^{2}(\Omega)}^{2}=\frac{1}{V^{2}}\sum_{|\mathbf{k}|>K_{\text{max}}}|\hat{\rho}(\mathbf{k})|^{2}\,|\hat{M}(\mathbf{k})|^{2}.$ (73)
Using Lemma 3,
$\frac{1}{V}\|\phi^{\mathrm{far}}-\phi_{t}^{\mathrm{far}}\|_{L^{2}(\Omega)}^{2}\leq\frac{C_{\rho}\|\bm{\rho}\|_{2}^{2}}{V^{2}}\sum_{|\mathbf{k}|>K_{\text{max}}}|\hat{M}(\mathbf{k})|^{2}.$ (74)
Since $|\hat{M}(\bm{k})|^{2}$ is radial and nonincreasing in $|\bm{k}|$, the lattice-to-continuum estimate of Lemma 6 (Appendix C.3) yields the bound
@@
$\sum_{|\bm{k}|>K_{\text{max}}}|\hat{M}(\bm{k})|^{2}\;\leq\;\Big(\frac{L}{2\pi}\Big)^{\!3}\int_{|\bm{\omega}|>\,\omega_{*}}|\hat{M}(\bm{\omega})|^{2}\,d\bm{\omega},\quad\omega_{*}=\max\!\Big\{\omega_{\max}-\tfrac{\sqrt{3}}{2}\tfrac{2\pi}{L},\,0\Big\}.$ (75)
Passing to spherical coordinates,
$\int_{|\bm{\omega}|>\omega_{*}}|\hat{M}(\bm{\omega})|^{2}\,d\bm{\omega}=4\pi\int_{\omega_{*}}^{\infty}|\hat{M}(\omega)|^{2}\,\omega^{2}\,d\omega.$ (76)
With $\hat{M}(\omega)=\tfrac{4\pi}{\omega^{2}}\hat{\gamma}_{\sigma}(\omega)$,
$\int_{\omega_{*}}^{\infty}|\hat{M}(\omega)|^{2}\,\omega^{2}\,d\omega=16\pi^{2}\int_{\omega_{*}}^{\infty}\frac{|\hat{\gamma}_{\sigma}(\omega)|^{2}}{\omega^{2}}\,d\omega.$ (77)
Collecting constants from Parseval’s theorem, the radial integral, and the lattice-to-continuum bound, and using $V=L^{3}$, we obtain the prefactor 8:
$\frac{1}{V^{2}}\sum_{|\bm{k}|>K_{\text{max}}}|\hat{M}(\bm{k})|^{2}\leq\frac{8}{V}\int_{\omega_{*}}^{\infty}\frac{|\hat{\gamma}_{\sigma}(\omega)|^{2}}{\omega^{2}}\,d\omega.$ (78)
Multiplying by $C_{\rho}\|\bm{\rho}\|_{2}^{2}$ completes the proof. ∎
Even though the PSWF function $\psi_{0}^{c_{s}}$ is implicit in nature, Theorem 1 allows us to evaluate its $L^{2}$ norm exactly for $\omega_{\max}=c_{s}/r_{c}$. The following proposition, which follows from Lemma 4, estimates the PSWF split-errors.
Proposition 1.
Let $\phi_{t}^{\mathrm{far}}$ denote the Fourier-space far-field sum truncated to modes $|k_{i}|\leq K_{\text{max}}$ for $i=1,2,3$, and let the PSWF mollifier be given by (54)–(55). Suppose that $\omega_{\max}=c_{s}/r_{c}$ and $\omega_{*}=\omega_{\max}-\tfrac{\sqrt{3}\pi}{L}$ so that $\omega_{*}\in(0,\omega_{\max})$, and define
$C_{*}\;:=\;1\;+\;\left(\frac{\omega_{\max}}{\omega_{*}}\right)^{\!2}\,\frac{c_{s}\,\lambda_{0}^{2}}{\,2\pi-c_{s}\,\lambda_{0}^{2}\,}.$ (79)
Then the squared RMS-error satisfies

### v2

@@ 3.2 Fast Ewald summation
The steps of the fast Ewald method, as presented above, are summarized in Algorithm 2 (that is also shown in a similar form in [19]).
@@ 4 Fast Ewald summation for the Laplace kernel using the first PSWF of order zero
In this section, we present how the first PSWF of order zero, $\psi_{0}^{c}$, can be employed both as a mollifier in the kernel split and as a window function for the spreading/interpolation steps of Algorithm 2.
4.1 General kernel split for numerical mollifiers with support [1,1][-1,1]
@@
Finally, substituting (55) in the kernel split $G=M+R$, the residual kernel becomes
@@
4.2 A PSWF mollifier for kernel splitting
Following Section 4.1, we now construct a PSWF mollifier. We start by defining a mollifier $\gamma^{c_{s}}_{1}$, supported on [1,1][-1,1], as
$\gamma_{1}^{c_{s}}(x):=\begin{cases}\dfrac{\psi_{0}^{c_{s}}(x)}{\lambda_{0}\psi_{0}^{c_{s}}(0)},&x\in[-1,1],\\[5.69054pt] 0,&x\notin[-1,1],\end{cases}$ (59)
where $\psi_{0}^{c_{s}}$ denotes the first PSWF with Fourier bandlimit $c_{s}>0$ (the subscript “s” stands for “split”). Here, $\lambda_{0}$ is the eigenvalue associated with $\psi_{0}^{c_{s}}$, and the normalization $\lambda_{0}\psi_{0}^{c_{s}}(0)$ follows from Lemma 2, ensuring that $\int_{\mathbb{R}}\gamma_{1}^{c_{s}}(x)\,\mathrm{d}x=1$.
To obtain a mollifier supported on a general radius $r_{c}>0$, we scale $\gamma^{c_{s}}_{1}$ as
@@ 4.2 A PSWF mollifier for kernel splitting
By Lemma 2, the Fourier transform of $\gamma_{r_{c}}^{c_{s}}$ can be written as
$\hat{\gamma}_{r_{c}}^{c_{s}}(\omega)=\hat{\gamma}_{1}^{c_{s}}(r_{c}\omega)=\frac{\psi_{0}^{c_{s}}\!\left(r_{c}\omega/c_{s}\right)}{\psi_{0}^{c_{s}}(0)}.$ (61)
For $|\omega|>c_{s}/r_{c}$, the argument $r_{c}\omega/c_{s}$ lies outside [1,1][-1,1], so this formula is no longer directly useful when $\psi_{0}^{c_{s}}$ is computed numerically from its representation on [1,1][-1,1]. In that case, the Fourier transform is instead evaluated from its defining integral:
$\hat{\gamma}_{r_{c}}^{c_{s}}(\omega)=\int_{-r_{c}}^{r_{c}}\gamma_{r_{c}}^{c_{s}}(x)\,e^{-\mathrm{i}\omega x}\,\mathrm{d}x=\frac{1}{\psi_{0}^{c_{s}}(0)}\int_{-1}^{1}\psi_{0}^{c_{s}}(u)e^{-i\omega r_{c}u}\,\mathrm{d}u.$ (62)
@@
$\Phi^{c_{s}}_{r_{c}}(x):=2\int_{0}^{x}\gamma^{c_{s}}_{r_{c}}(t)\,\mathrm{d}t=\frac{2}{\lambda_{0}\,\psi_{0}^{c_{s}}(0)}\int_{0}^{x/r_{c}}\psi_{0}^{c_{s}}(u)\,\mathrm{d}u,\qquad x\in[0,r_{c}],$ (63)
@@
$\hat{M}(\bm{\omega})=\frac{4\pi}{|\bm{\omega}|^{2}}\,\hat{\gamma}_{r_{c}}^{c_{s}}(|\bm{\omega}|),$ (64)
@@
Table 1: Gaussian and PSWF-based mollifiers. The Gaussian mollifier has width parameter $\sigma$, while the PSWF mollifier is compactly supported with cutoff $r_{c}$ and bandlimit $c_{s}$.
@@
$\gamma_{r_{c}}^{\text{PSWF}}(x)=\begin{cases}\dfrac{\psi_{0}^{c_{s}}(x/r_{c})}{r_{c}\,\lambda_{0}\,\psi_{0}^{c_{s}}(0)},&x\in[-r_{c},r_{c}]\\[5.81247pt] 0,&\text{otherwise}\end{cases}$ $\hat{\gamma}_{r_{c}}^{\text{PSWF}}(\omega)=\displaystyle\frac{\psi_{0}^{c_{s}}\!\left(r_{c}\omega/c_{s}\right)}{\psi_{0}^{c_{s}}(0)}$
@@ 4.3 A PSWF window function
We define a one-dimensional window function by
$\varphi_{1}^{c_{w}}(x):=\begin{cases}\displaystyle\frac{\psi_{0}^{c_{w}}(x)}{\psi_{0}^{c_{w}}(0)},&x\in[-1,1],\\[5.0pt] 0,&\text{otherwise}.\end{cases}$ (65)
By Lemma 2, its Fourier transform is given by
@@
$\varphi^{c_{w}}(\bm{x}):=\prod_{i=1}^{3}\varphi_{1}^{c_{w}}(x_{i}/\alpha)=\begin{cases}\displaystyle\prod_{i=1}^{3}\frac{\psi_{0}^{c_{w}}(x_{i}/\alpha)}{\psi_{0}^{c_{w}}(0)},&\bm{x}\in[-\alpha,\alpha]^{3},\\[5.0pt] 0,&\text{otherwise}.\end{cases}$ (67)
Since the Fourier transform of a tensor product factorizes into the product of the one-dimensional transforms, we obtain
@@
The PSWF window function $\varphi^{c_{w}}$ is by definition compactly supported on $\alpha,\alpha]^{3}$, hence it equals its truncation, $\varphi_{t}^{c_{w}}(\bm{x})=\varphi^{c_{w}}(\bm{x})$.
@@
Thus, for force computations (e.g. $\bm{F}_{i}=-\rho_{i}\nabla\phi_{i}$), $\nabla\phi_{i}$ can be obtained by replacing the last interpolation step (step 5) in Algorithm 2 with
@@ 4.3.1 Comparison of the PSWF window with Gaussian and B-spline windows
$\varphi^{\text{PSWF}}(x)=\begin{cases}\dfrac{\psi_{0}^{c_{w}}(x/\alpha)}{\psi_{0}^{c_{w}}(0)},&x\in[-\alpha,\alpha]\\[5.81247pt] 0,&\text{otherwise}\end{cases}$ $\hat{\varphi}^{\text{PSWF}}(\omega)=\displaystyle\alpha\,\lambda_{0}\,\frac{\psi_{0}^{c_{w}}(\alpha\omega/c_{w})}{\psi_{0}^{c_{w}}(0)}$
@@
B-splines have support width $P$, equal to the spline order, and have Fourier transform $\hat{\varphi}(\omega)=h\,\hat{B}_{P}(h\omega)$, where $h$ denotes the grid spacing and $\hat{B}_{P}(\omega)=(\sin(\omega/2)/(\omega/2))^{P}$. Consequently, their spectral decay is algebraic. For a fixed real-space width $\alpha$, B-splines therefore require either a finer grid or a higher order $P$ to match the accuracy of the PSWF window; increasing $P$ enlarges the support and increases the cost of spreading and interpolation. The widely used SPME method [11]—implemented, for example, in GROMACS [35]—employs non-centered cardinal B-splines for spreading and interpolation and deconvolves their spectrum through a modified diagonal scaling (see Remark 5). However, the underlying algebraic spectral decay remains unchanged.
@@ 5 Theoretical error analysis for the PSWF split and window functions
In this section, we present a rigorous theoretical analysis of the error sources arising in fast Ewald summation using PSWF-based functions. We derive absolute $L^{2}$ error estimates associated with both the PSWF mollifier and the PSWF window function introduced in Section 4. For the purpose of parameter selection—where fully rigorous bounds are unnecessarily detailed—we also derive simplified, closed-form error models that are well suited for practical use.
@@
Let $\phi^{\text{far}}_{t}$ denote the truncated far-field sum $\phi^{\text{far}}$ (see (43)), and let $\phi^{\text{far}}_{t,h}$ denote an approximation of $\phi^{\text{far}}_{t}$ computed using the fast Ewald algorithm (Algorithm 2). In fast Ewald summation, two principal sources of error arise: (i) truncation of Fourier modes in the Fourier-transformed mollified kernel—referred to as split errors—quantified by $\frac{1}{\sqrt{V}}\|\phi^{\text{far}}-\phi_{t}^{\text{far}}\|_{L^{2}(\Omega)}$, and (ii) approximation errors arising in the spreading and interpolation steps, quantified by $\frac{1}{\sqrt{V}}\|\phi_{t}^{\text{far}}-\phi_{t,h}^{\text{far}}\|_{L^{2}(\Omega)}$.
@@
5.1 Kernel split truncation errors
@@
The constant $C_{\rho}$ reflects the degree of cancellation among the oscillatory exponential terms. In the worst case (highly clustered particles), one may have $C_{\rho}$ as large as $n$. For well-distributed configurations (e.g., approximately uniform particle distributions), the exponential vectors corresponding to distinct wavenumbers are nearly orthogonal, leading to substantial cancellations. In such cases $C_{\rho}\approx 1$.
The following proposition provides an asymptotic bound on the Fourier-tail error induced by the PSWF kernel split.
Proposition 1.
Assume that $c_{s}\geq\sqrt{3}$. Then for $r_{c}/c_{s}$ sufficiently small,
$\frac{1}{V}\,\|\phi^{\mathrm{far}}-\phi_{t}^{\mathrm{far}}\|_{L^{2}(\Omega)}^{2}\leq\frac{8C_{\rho}\,\|\bm{\rho}\|_{2}^{2}\,r_{c}}{3Vc_{s}}\left(\frac{\psi_{0}^{c_{s}}(1)}{\psi_{0}^{c_{s}}(0)}\right)^{2}\left(1+\mathcal{O}\!\left(\frac{r_{c}}{c_{s}}\right)\right),$ (77)
where $1\leq C_{\rho}\leq n$ is the concentration constant from Lemma 6.
@@
$\frac{1}{V}\|\phi^{\mathrm{far}}-\phi_{t}^{\mathrm{far}}\|_{L^{2}(\Omega)}^{2}$ $=\frac{1}{V^{2}}\sum_{\bm{k}\notin\mathcal{I}_{\bm{m}}}|\hat{\rho}(\bm{k})|^{2}\,|\hat{M}(\bm{\omega}(\bm{k}))|^{2}$
$\leq\frac{1}{V^{2}}\sum_{|\bm{k}|>K_{\max}}|\hat{\rho}(\bm{k})|^{2}\,|\hat{M}(\bm{\omega}(\bm{k}))|^{2}.$ (78)
Applying Lemma 6 yields
$\frac{1}{V}\|\phi^{\mathrm{far}}-\phi_{t}^{\mathrm{far}}\|_{L^{2}(\Omega)}^{2}\leq\frac{C_{\rho}\|\bm{\rho}\|_{2}^{2}}{V^{2}}\sum_{|\bm{k}|>K_{\text{max}}}|\hat{M}(\bm{\omega}(\bm{k}))|^{2}.$ (79)
It therefore remains to estimate the tail sum.
Recall that
$\hat{M}(\bm{\omega}(\bm{k}))=\frac{4\pi}{|\bm{\omega}(\bm{k})|^{2}}\frac{\psi_{0}^{c_{s}}(r_{c}|\bm{\omega}(\bm{k})|/c_{s})}{\psi_{0}^{c_{s}}(0)}.$ (80)
For $|\bm{k}|>K_{\text{max}}$ we have $|\bm{\omega}(\bm{k})|>\omega_{\max}=c_{s}/r_{c}$, and hence $r_{c}|\bm{\omega}(\bm{k})|/c_{s}>1$. Therefore, by Lemma 4, which gives $|\psi_{0}^{c_{s}}(s)|\leq\psi_{0}^{c_{s}}(1)/s$ for $s\geq 1$ and $c_{s}\geq\sqrt{3}$, we obtain
$|\hat{M}(\bm{\omega}(\bm{k}))|\leq\frac{4\pi c_{s}}{r_{c}}\frac{\psi_{0}^{c_{s}}(1)}{\psi_{0}^{c_{s}}(0)}\frac{1}{|\bm{\omega}(\bm{k})|^{3}}.$ (81)
Squaring gives
$|\hat{M}(\bm{\omega}(\bm{k}))|^{2}\leq\frac{16\pi^{2}c_{s}^{2}}{r_{c}^{2}}\left(\frac{\psi_{0}^{c_{s}}(1)}{\psi_{0}^{c_{s}}(0)}\right)^{2}\frac{1}{|\bm{\omega}(\bm{k})|^{6}}.$ (82)
Summing over the lattice points $\bm{\omega}(\bm{k})=(2\pi/L)\bm{k}$ with $|\bm{k}|>K_{\text{max}}$ and estimating the resulting discrete sum by a radial integral via Lemma 8, we obtain
@@
$\sum_{|\bm{k}|>K_{\text{max}}}|\hat{M}(\bm{\omega}(\bm{k}))|^{2}$ $\leq\frac{16\pi^{2}c_{s}^{2}}{r_{c}^{2}}\left(\frac{\psi_{0}^{c_{s}}(1)}{\psi_{0}^{c_{s}}(0)}\right)^{2}\left(\frac{L}{2\pi}\right)^{3}\int_{|\bm{\omega}|>\omega_{\max}-\tfrac{\sqrt{3}\pi}{L}}\frac{1}{|\bm{\omega}|^{6}}\,\mathrm{d}\bm{\omega}$